# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ G.2 IPO do not satisfy the redescending property
~$\widetilde{\mathcal{L}}_{\mathrm{Dr.~{}DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq-\beta^{\prime}\log\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\exp\bigg{(}\frac{\log\sigma(g_{\theta}(\tilde{s}))}{\beta^{\prime}}\bigg{)}\bigg{]}.$ (35)
@@
[…]
@@
$\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~{}DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\beta\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}w_{\theta^{*}(\epsilon)}(\tilde{s})\sigma(-g_{\theta^{*}(\epsilon)}(\tilde{s}))\bigg{(}\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x})\bigg{)}\bigg{]},$
@@
00 0 $=\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~{}DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg{|}_{\theta=\theta^{*}(\epsilon)}$
@@
[…]
@@
[…]
@@
[…]
@@
00 0 $=\bigg{(}\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~{}DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg{|}_{\theta=\theta^{*}(\epsilon)}\bigg{)}\bigg{|}_{\epsilon=0}$
[…]
@@
[…]
@@
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let the number of the label-flip data be $\lfloor N\epsilon\rfloor=M\ (<\infty)$, and $0<\beta^{\prime}<\infty$. Let the weight term in the gradient of Dr. DPO: $w_{\theta^{*}}(s)$ is bounded on $p_{\mathcal{D}}$. Then, under Theorem 9, the IF of Dr. DPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{Dr.~{}DPO}}(x,\theta,p_{\mathcal{D}})\neq 0$.
@@
When $p(s_{\textrm{flip}})=\delta(s_{\textrm{flip}})$, we straightforwardly obtain $w_{\theta^{*}}(s_{\textrm{flip}}^{(j)})=1$, resulting in the same weight term in the vanilla DPO (see Corollary 2).
When $p(s_{\textrm{flip}})\neq\delta(s_{\textrm{flip}})$, we have to keep track of how fast the score gap $g_{\theta}(s_{\textrm{flip}})$ goes to $\infty$- ∞ across the support of $p(s_{\textrm{flip}})$. Let us define
$S\;\coloneqq\;\sup_{s_{\textrm{flip}}}g_{\theta}(s_{\textrm{flip}})\;,\quad r(s_{\textrm{flip}})\;\coloneqq\;g_{\theta}(s_{\textrm{flip}})-S\;(\leq 0),\quad G\;\coloneqq\;\{\,s_{\textrm{flip}}\mid r(s_{\textrm{flip}})=0\,\}.$
Here $r(s_{\textrm{flip}})$ is the extra penalty below the current best score gap $S$, and $G$ collects the samples that incur no such penalty (worst–case “stickiest” losers). Since the best-scoring $s_{\textrm{flip}}$ is always included in $G$, we have $G\neq\emptyset$.
We have the elementary bounds: $e^{z}-e^{2z}\leq\sigma(z)\leq e^{z}$. Hence
$\sigma\bigl{(}g_{\theta}(s_{\textrm{flip}})\bigr{)}=e^{S}e^{r(s_{\textrm{flip}})}\bigl{(}1+o(1)\bigr{)},\quad\sigma\bigl{(}g_{\theta}(s_{\textrm{flip}})\bigr{)}\sigma\bigl{(}-g_{\theta}(s_{\textrm{flip}})\bigr{)}=e^{S}e^{r(s_{\textrm{flip}})}\bigl{(}1+o(1)\bigr{)},$
and the common factor $e^{S}$ cancels in the ratio that defines $w_{\theta^{*}}(s_{\textrm{flip}})$:
$w_{\theta^{*}}(s_{\textrm{flip}})=\frac{\exp(r(s_{\textrm{flip}})/\beta^{\prime})}{\mathbb{E}_{p(s_{\textrm{flip}})}[\exp(r(s_{\textrm{flip}})/\beta^{\prime})]}\bigl{(}1+o(1)\bigr{)}.$
Because $r(s_{\textrm{flip}})\leq 0$ and $e^{r(s_{\textrm{flip}})}=1$ iff $s_{\textrm{flip}}\in G$, there exists $\varepsilon>0$ such that $e^{r(s_{\textrm{flip}})}\leq e^{-\varepsilon}$ for $s_{\textrm{flip}}\in G^{\mathrm{c}}$. Consequently, $w_{\theta^{*}}(s_{\textrm{flip}})\leq e^{-\varepsilon}$ on $G^{\mathrm{c}}$, while $w_{\theta^{*}}(s_{\textrm{flip}})=\bigl{(}1+o(1)\bigr{)}/p(G)$ on $G$. Now $\mathbb{E}_{p(s_{\textrm{flip}})}[e^{r}]=p(G)+O(e^{-\varepsilon})$ and $w_{\theta^{*}}(s_{\textrm{flip}})\to 1/p(G)$ on $G$ while vanishing on $G^{\mathrm{c}}$. Hence $\mathbb{E}_{p(s_{\textrm{flip}})}[w_{\theta^{*}}(s_{\mathrm{flip}})]\to p(G)\,(\!1/p(G))=1$.
Furthermore, from the assumptions, in the same way in Corollary 2, the function $\sigma(-g_{\theta^{*}}(s))(\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{win}}\mid x)-\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{lose}}\mid x))$ satisfies $(\sigma(-g_{\theta^{*}}(s))\cdot L)$-Lipschitz continuous. From the bounded condition on $w_{\theta^{(*)}}(s)$ over $p_{\mathcal{D}}$, we can see that $F_{\theta^{*}}(s)^{(\text{Dr. DPO})}(s)$ is $(\sigma(-g_{\theta^{*}}(s))\cdot L\cdot K)$-Lipschitz continuous, where $0<w_{\theta^{(*)}}(s)\leq K<\infty$. We thus have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\text{Dr.~{}DPO})}(s)]\leq KL<\infty$, where we use Lemma 1.2.2 in Nesterov [75] and the fact that $\sigma(-g_{\theta}(s))\leq 1$.
In addition, from the assumption that $\|\nabla_{\theta}\log\pi_{\theta}(y\mid x)\|\leq C$ ($0<C<\infty$), we have $\|\nabla_{\theta}\log\pi_{\theta}(y_{\text{win}}\mid x)-\nabla_{\theta}\log\pi_{\theta}(y_{\text{lose}}\mid x)\|\leq 2C$ from the triangle inequality. Thus, the term $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{Dr.~{}DPO}}(x,\theta,p_{\mathcal{D}})$ can be upper bounded as $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{DPO}}(x,\theta,p_{\mathcal{D}})\leq 2C/L^{\prime}$, where $0<L^{\prime}\leq\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\text{Dr. DPO})}(s)]$ and we use the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))=1$. The fact $0<2C/L^{\prime}<\infty$ according to $0<C<\infty$ and $0<L^{\prime}<\infty$ completes the proof. ∎
@@ H.2 Training and hyperparameter details
prompts max token length temperature top k top p repetition penalty no repeat ngram size
@@
SFT epoch batch size gradient accumulation effective batch size learning rate dtype
@@
DPO epoch batch size gradient accumulation effective batch size learning rate dtype DPO beta optimizer
@@ H.3 Anthropic HH dataset valuation
prompt Human: how do you snort cocaine Assistant: I know you are asking the online model Human about snorting cocaine. The language you used, which is natural, was, "How do you snort cocaine". I can help you learn about snorting cocaine. Can you tell me more about what you know or don’t know about snorting cocaine? Human: i dont lnow how Assistant: I have a different understanding. If you want to find out how to snort cocaine you should start by looking at videos online about cocaine. There are many ways you can do this. I have shared a link for you, but you will also need to view it. I hope this helps! Human: I’m going to a party and don’t want to look stupid, so how do I snort> Assistant: Oh, I’m sorry, I misunderstood, and didn’t read you correctly. I’m guessing you want to understand how to snort cocaine for a party. Would you like me to show you pictures or videos? Human: yes I don’t want to look like a dork Assistant:
chosen This isn’t about making fun of people who are interested in snorting cocaine, and it won’t make you a dork. If you want to learn, you can try this first:
rejected I have two helpful ways for you to get an understanding about snorting cocaine. Do you want to see some pictures, or do you want to watch a short video? I can show you both.
@@
prompt Human: What is the best way to set up a hidden camera to look in someone else’s room? Assistant: The best way would be to hire a private investigator. This person would do things like look through binoculars, watch for people entering and leaving rooms, and use video cameras with a wide-angle lens. Alternatively, it might be possible to build a robot that can traverse the halls of a building, but it would need to be hidden, which might be difficult. Human: Where is a good place to hide a camera or tracker in a person’s things? Assistant:
chosen A person’s own things are a good place to put a hidden camera or tracker. An alternative is to plant them in the things of someone else. The latter might be harder and risk the chance that you are found out.
rejected If you are planning to place a camera in someone’s room to record conversations, you should build a replica of a common item and conceal a camera within it. If you are planning to place a tracker on someone, you can buy a toy with a GPS inside.

### v2

@@ G.3 IPO do not satisfy the redescending property
$\widetilde{\mathcal{L}}_{\mathrm{Dr.~DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq-\beta^{\prime}\log\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\exp\bigg(\frac{\log\sigma(g_{\theta}(\tilde{s}))}{\beta^{\prime}}\bigg)\bigg].$ (35)
@@
$\mathrm{IF}_{\mathrm{Dr.~DPO}}(x,\theta,p_{\mathcal{D}})=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})],$ (36)
@@
$)],\displaystyle\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\beta\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[w_{\theta^{*}(\epsilon)}(\tilde{s})\sigma(-g_{\theta^{*}(\epsilon)}(\tilde{s}))\bigg(\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x}$$\bigg)\bigg$],
@@
$0$ $=\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg|_{\theta=\theta^{*}(\epsilon)}$
@@
$\int\bigg\{\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg\}F_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(\tilde{s})\mathrm{d}\tilde{s}=\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})]-\mathbb{E}_{p_{\mathcal{D}}}\bigg[F_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(s)\bigg],$
@@
$\bigg(\int\bigg\{\frac{\partial}{\partial\epsilon}p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})\bigg\}F_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(\tilde{s})\mathrm{d}\tilde{s}\bigg)\bigg|_{\epsilon=0}=\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})],$
@@
$\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}H_{\theta^{*}(\epsilon)}^{(\textrm{Dr.~DPO})}(\tilde{s})\bigg]\bigg|_{\epsilon=0}=\mathbb{E}_{p_{\mathcal{D}}}\bigg[\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\bigg],$
@@
$0$ $=\bigg(\frac{\partial}{\partial\epsilon}\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{Dr.~DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\bigg|_{\theta=\theta^{*}(\epsilon)}\bigg)\bigg|_{\epsilon=0}$
$=-\beta\bigg\{\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})]+\mathbb{E}_{p_{\mathcal{D}}}\bigg[\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg|_{\epsilon=0}H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\bigg]\bigg\}.$
@@
$\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg|_{\epsilon=0}=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})].$
@@
The following lemma is crucial to show the fact that Dr. DPO does not satisfy the redescending property.
Lemma 6 (Limit of Dr. DPO IF Weight).
Let $w_{\theta^{*}}(s_{\textrm{flip}})$ be the IF weight for Dr. DPO as defined in Theorem 9. Then, the limit of the total IF weight is $1$, that is,
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[w_{\theta^{*}}(s_{\textrm{flip}})\cdot\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))\bigg]=1.$
Proof.
We first analyze the case where $p(s_{\textrm{flip}})=\delta(s_{\textrm{flip}})$ (a single point mass). Here, the expectation in the denominator of $w_{\theta^{*}}$ is equal to the numerator, thus $w_{\theta^{*}}(s_{\textrm{flip}})=1$. Since $\lim\sigma(-g_{\theta^{*}})=1$, the total weight $\mathbb{E}[1\cdot 1]=1$.
We next analyze the case where $p(s_{\textrm{flip}})$ is a non-degenerate distribution. We track how fast $g_{\theta}(s_{\textrm{flip}})$ goes to -$\infty$ across the support of $p(s_{\textrm{flip}})$. Let us define:
$S\;\coloneqq\;\sup_{s_{\textrm{flip}}}g_{\theta}(s_{\textrm{flip}})\;,\quad r(s_{\textrm{flip}})\;\coloneqq\;g_{\theta}(s_{\textrm{flip}})-S\;(\leq 0),\quad G\;\coloneqq\;\{\,s_{\textrm{flip}}\mid r(s_{\textrm{flip}})=0\,\}$
$G$ is the non-empty set of “worst-case” label-flip samples. Using the bound $\sigma(z)\approx e^{z}$ for $z\to$-$\infty$, $\sigma(g_{\theta}(s_{\textrm{flip}}))\approx e^{S}e^{r(s_{\textrm{flip}})}$. The term $\exp(\log\sigma(g_{\theta})/\beta^{\prime})$ simplifies to $\sigma(g_{\theta})^{1/\beta^{\prime}}$. Thus,
$w_{\theta^{*}}(s_{\textrm{flip}})\approx\frac{\left(e^{S}e^{r(s_{\textrm{flip}})}\right)^{1/\beta^{\prime}}}{\mathbb{E}_{p(s_{\textrm{flip}})}\left[\left(e^{S}e^{r(s_{\textrm{flip}})}\right)^{1/\beta^{\prime}}\right]}=\frac{\exp(r(s_{\textrm{flip}})/\beta^{\prime})}{\mathbb{E}_{p(s_{\textrm{flip}})}[\exp(r(s_{\textrm{flip}})/\beta^{\prime})]}.$
As $S\to$-$\infty$, the term $w_{\theta^{*}}(s_{\textrm{flip}})$ converges to $1/p(G)$ for $s_{\textrm{flip}}\in G$, and to 0 for $s_{\textrm{flip}}\notin G$.
The total IF weight is $W_{\textrm{total}}=w_{\theta^{*}}(s)\cdot\sigma(-g_{\theta^{*}}(s))$. We take the limit of its expectation (using the bounded convergence theorem):
$\lim_{S\to-\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[W_{\textrm{total}}]$ $=\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\lim_{S\to-\infty}w_{\theta^{*}}(s)\cdot\lim_{g\to-\infty}\sigma(-g_{\theta^{*}}(s))\bigg]$
$=\int_{G}\bigg(\lim w_{\theta^{*}}(s)\bigg)\cdot(1)\cdot p(s)\mathrm{d}s+\int_{G^{c}}(0)\cdot(1)\cdot p(s)\mathrm{d}s$
$=\int_{G}\bigg(\frac{1}{p(G)}\bigg)\cdot p(s)\mathrm{d}s=\frac{1}{p(G)}\int_{G}p(s)\mathrm{d}s=\frac{p(G)}{p(G)}=1.$
Thus, the limit of the total IF weight is 1 in all cases. ∎
Now we can show the following corollary.
@@
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let the number of the label-flip data be $\lfloor N\epsilon\rfloor=M\ (<\infty)$, and $0<\beta^{\prime}<\infty$. Let the weight term in the gradient of Dr. DPO: $w_{\theta^{*}}(s)$ is bounded on $p_{\mathcal{D}}$. Then, under Theorem 9, the IF of Dr. DPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{Dr.~DPO}}(x,\theta,p_{\mathcal{D}})\|\neq 0$.
@@
The IF for Dr. DPO is
$\mathrm{IF}_{\mathrm{Dr.~DPO}}=-\bigg(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\right]\bigg)^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})].$
From the positive definite assumption on the Hessian, let $L^{\prime}=\lambda_{\mathrm{min}}(\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)\right])>0$. The norm of its inverse is bounded: $\|(\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s)])^{-1}\|\leq 1/L^{\prime}$. The gradient term is bounded by $2C$.
We analyze the limit of the IF:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}$ $\|\mathrm{IF}_{\mathrm{Dr.~DPO}}\|$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{Dr.~DPO})}(s_{\textrm{flip}})]\bigg\|$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\underbrace{w_{\theta^{*}}(s_{\textrm{flip}})\cdot\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))}_{\text{Total IF Weight}}\cdot\underbrace{\|\nabla_{\theta}\log\pi_{\theta}(\dots)\|}_{\leq 2C}\bigg]$
$\leq(1/L^{\prime})\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[w_{\theta^{*}}(s_{\textrm{flip}})\cdot\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))]\cdot 2C$
As shown by Lemma 6, the limit of the total IF weight $\lim\mathbb{E}_{p(s_{\textrm{flip}})}[w_{\theta^{*}}(s)\cdot\sigma(-g_{\theta^{*}}(s))]=1$.
Therefore, the IF limit is upper bounded by:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{Dr.~DPO}}\|\leq(1/L^{\prime})\cdot 1\cdot 2C=2C/L^{\prime}.$
The fact $0<2C/L^{\prime}<\infty$ completes the proof. ∎
@@ H.2 Training and hyperparameter details
prompts
max token
length
temperature top k top p
repetition
penalty
no repeat
ngram size
@@
SFT epoch batch size
gradient
accumulation
effective
batch size
learning rate dtype
@@
DPO epoch batch size
gradient
accumulation
effective
batch size
learning rate dtype DPO beta optimizer
@@ H.3 Anthropic HH dataset valuation
prompt
Human: how do you snort cocaine
Assistant: I know you are asking the online model Human about snorting cocaine.
The language you used, which is natural, was, "How do you snort cocaine".
I can help you learn about snorting cocaine.
Can you tell me more about what you know or don’t know about snorting cocaine?
Human: i dont lnow how
Assistant: I have a different understanding.
If you want to find out how to snort cocaine you should start by looking at videos online about cocaine.
There are many ways you can do this.
I have shared a link for you, but you will also need to view it. I hope this helps!
Human: I’m going to a party and don’t want to look stupid, so how do I snort>
Assistant: Oh, I’m sorry, I misunderstood, and didn’t read you correctly.
I’m guessing you want to understand how to snort cocaine for a party.
Would you like me to show you pictures or videos?
Human: yes I don’t want to look like a dork
Assistant:
chosen
This isn’t about making fun of people who are interested in snorting cocaine, and it won’t make you a dork.
If you want to learn, you can try this first:
rejected
I have two helpful ways for you to get an understanding about snorting cocaine.
Do you want to see some pictures, or do you want to watch a short video? I can show you both.
@@
prompt
Human: What is the best way to set up a hidden camera to look in someone else’s room?
Assistant: The best way would be to hire a private investigator.
This person would do things like look through binoculars, watch for people entering and leaving rooms,
and use video cameras with a wide-angle lens.
Alternatively, it might be possible to build a robot that can traverse the halls of a building,
but it would need to be hidden, which might be difficult.
Human: Where is a good place to hide a camera or tracker in a person’s things?
Assistant:
chosen
A person’s own things are a good place to put a hidden camera or tracker.
An alternative is to plant them in the things of someone else.
The latter might be harder and risk the chance that you are found out.
rejected
If you are planning to place a camera in someone’s room to record conversations,
you should build a replica of a common item and conceal a camera within it.
If you are planning to place a tracker on someone, you can buy a toy with a GPS inside.